OpenStudy (anonymous):

any math student or teacher here?plzsolve this question,,if a+b is not 0,than the eqation a(x-a)=2ab-b(x-b) has a solution x=a+b

OpenStudy (anonymous):

plzzz review yours second step...a(x-a)+(bx-b)-2ab... its should be ax-a2+bx-b-2ab not b2.... plz clear

OpenStudy (anonymous):

\[a(x−a)+b(x−b)−2ab=0\] \[ax−a^2+bx−b^2−2ab=0\] \[x(a+b)=(a^2+b^2+2ab)\] Use: \[a^2+b^2+2ab=(a+b)^2\] \[x \cancel{(a+b)}= \cancel{(a+b)} \times (a+b)\] \[x=(a+b)\] which is the required solution.

OpenStudy (anonymous):

yes now tyou are right........gud ,,weldone i m feeling happy coz tommorow i will hv submit answer ...thanx a lot